Exercise 14.58 - Enhanced with Feedback 26 of 48 Review I Constants I Periodic Table The half-life for the radioactive decayThe half-life for the radioactive decay of C-14 is 5730 years. You may want to reference (Pages 598 - 605) Section 14.5 while completing this problem.
Answer #1

All radioactive decay processes follows first order kinetics.

Fisust ostdesi ate constant is given as, t (a-n)a = initial concentration.

a-x = concentration after time 't'

Half-lite is given as, ts balar 0.693 tve k thy =5730 ys a= loo% PR (a- = 100-30 = 70% k=0.693 5730 S88 X10 Dreep/crotx01e =o